-1(-6x-5)-3 = 2(3x+4)-2
Answer by Edwin McCravy(20055)

-1(-6x-5)-3 = 2(3x+4)-2
                               Distribute the -1 on the left and the 2 on the right

     6x+5-3 = 6x+8-2

       6x+2 = 6x+6
      -6x    -6x 
---------------------
          2 =    6

2 can never be equal to 6 no matter what numerical value
the variable x takes on, soo there is no solution.

The solution set is Ø, the empty set.

Edwin